In a deck of 52 card If you draw 3 cards and replace after each draw what is the probability of picking hearts one or more times?
I think it might be 12/52.. there are 12 hearts 52 cards and since its being replaced the probabilty will stay the same? OR is it 12/52+12/52+12/52=36/52?

Answer by stanbon(75887)

In a deck of 52 card If you draw 3 cards and replace after each draw what is the probability of picking hearts one or more times?
I think it might be 12/52.. there are 12 hearts 52 cards and since its being replaced the probabilty will stay the same? OR is it 12/52+12/52+12/52=36/52?
------------------
Since you are replacing the card after drawing, the events are independent.
This is a binomial problem with n=3, p=1/4, 1<=x<=3
---------------
P(at least one heart) = 1 - P(no hearts)
= 1 - (3/4)^3 = 0.5781
=========================
Cheers,
Stan H.